walt made an extra $5000 last year from a part-time job. He invested part of the money at 9% and the rest at 7%. He made a total of $370 in interest. how much was invested at 7%?
Answer by checkley77(12569)

.09x+.07(5,000-x)=370
.09x+350-.07x=370
.02x=370-350
.02x=20
x=20/.02
x=$1,000 invested @ 9%.
5,000-100=4,000 invested @ 7%
Proof:
.09*1,000+.07*4,000=370
90+280=370
370=370